OpenStudy (anonymous):

LCD of 10/11 and 56/121

OpenStudy (campbell_st):

try 121

OpenStudy (mokeira):

 that is GCD @campbell_st

OpenStudy (anonymous):

110/121 and 56/121

OpenStudy (anonymous):

I checked it with online LCD calculator: 10/11 = 110/121 56/121 = 56/121 http://www.hackmath.net/en/calculator/least-common-denominator?input=10%2F11++56%2F121&submit=Calculate+LCD